# AnimateAnywhere: Rouse the Background in Human Image Animation https://animateanywhere.github.io/

### v1

@@
Abstract.
Human image animation aims to generate human videos of given characters and backgrounds that adhere to the desired pose sequence. However, existing methods focus more on human actions while neglecting the generation of background, which typically leads to static results or inharmonious movements. The community has explored camera pose-guided animation tasks, yet preparing the camera trajectory is impractical for most entertainment applications and ordinary users. As a remedy, we present an AnimateAnywhere framework, rousing the background in human image animation without requirements on camera trajectories. In particular, based on our key insight that the movement of the human body often reflects the motion of the background, we introduce a background motion learner (BML) to learn background motions from human pose sequences. To encourage the model to learn more accurate cross-frame correspondences, we further deploy an epipolar constraint on the 3D attention map. Specifically, the mask used to suppress geometrically unreasonable attention is carefully constructed by combining an epipolar mask and the current 3D attention map. Extensive experiments demonstrate that our AnimateAnywhere effectively learns the background motion from human pose sequences, achieving state-of-the-art performance in generating human animation results with vivid and realistic backgrounds. The source code and model will be available at https://github.com/liuxiaoyu1104/AnimateAnywhere.
1. Introduction
@@
Recent advancements in generative diffusion models have significantly advanced the field of image animation (Hu, 2024; Xu et al., 2024b; Zhu et al., 2024; Zhou et al., 2024; Wang et al., 2024e, d), enabling the creation of highly realistic and controllable human image animation, particularly in terms of identity preservation and motion consistency. Despite these advancements, most existing human image animation methods focus on improving the generation quality of foreground humans while leaving the background static or unconstrained. However, in practice, the photographers usually need to adjust the cameras to better capture human movements, postures, and etc., which results in background changes. Thus, the static or disharmonious backgrounds are insufficient to meet the requirements for realistic and diverse human video generation.
To address this issue, one possible solution is to employ control signals to guide background motion. For example, Humanvid (Wang et al., 2024a) controls the background motion via camera trajectories estimated from the reference videos, while Liu et al. (Liu et al., 2024) leverage sparse tracking points and human pose sequences to independently control background and foreground motions. However, these methods typically acquire background control signals from a reference video, which is unavailable in many scenarios. Although camera trajectories can be manually designed, this easily results in disharmony between human and background motion. It requires professional skills and is redundant for most applications and ordinary users. Additionally, another limitation of these methods is that a reference image with both the human and the background is required.
@@
Specifically, we build the AnimateAnywhere framework upon a pre-trained image-to-video model (i.e., CogVideoX (Yang et al., 2024b)) and independently inject a reference human image, a reference background image, and a human pose sequence into the generation process. Thus, AnimateAnywhere supports free combinations of reference humans, backgrounds, and pose sequences, enabling highly flexible and controllable video generation. To realize the key idea of AnimateAnywhere, we propose the Background Motion Learner (BML) module, which learns background motion from the latent features of human pose sequences by applying LoRA (Hu et al., 2022) on the 3D-attention-based block. Moreover, although 3D full attention allows each background pixel to attend to any other pixel from all frames, this highly free attention manner also brings difficulties to optimization, easily resulting in geometrically inconsistent background generation. To further improve the cross-frame geometric consistency of the background, we propose an adaptive epipolar constraint for the 3D attention map. Therein, a mask used for suppressing geometrically unreasonable attention is elaborately designed from an epipolar mask and the current 3D attention map.
@@
We propose a Background Motion Learner (BML) module to predict background motion from human pose sequences and an adaptive epipolar constraint for the 3D attention map to enhance background geometric consistency.
@@
2. Related Work
2.1. Human Image Animation
@@
Human image Animation aims to generate coherent human videos from a reference image and a target pose sequence. By leveraging the generative visual capabilities of diffusion models, recent advancements (Hu, 2024; Xu et al., 2024b; Zhu et al., 2024; Zhou et al., 2024; Wang et al., 2024e, d, a; Yao et al., 2025) have enabled the generation of high-quality and consistent human animation videos. AnimateAnyone (Hu, 2024) and MagicAnimate (Xu et al., 2024b) introduce a video diffusion model and a reference network for temporal consistency and identity preservation, respectively. Champ (Zhu et al., 2024) incorporates four distinct control signals to provide comprehensive 3D shape and detailed pose guidance. Following these, some works focus on realistic hands (Zhou et al., 2024; Wang et al., 2024e), efficient temporal modeling (Wang et al., 2024d), or a pose alignment algorithm (Tong et al., 2024) to further improve the performance of human animation. Animate-X (Tan et al., 2024) extends animation to diverse character types with a unified framework.
@@
Despite achieving reasonable results in maintaining temporal consistency in humans, these approaches face challenges in generating realistic and dynamic backgrounds. MIMO (Men et al., 2024) and AnimateAnyone2 (Hu et al., 2025) address this challenge by taking the background video directly from a reference video as input. Liu et al. (Liu et al., 2024) extract sparse point tracking from the reference video and use it to guide background animation. Humanvid (Wang et al., 2024a) utilizes a SLAM-based method (Wang et al., 2024b; Teed and Deng, 2021) to estimate an accurate camera trajectory from the original videos and leverages CameraCtrl (He et al., 2024) and Animate Anyone (Hu, 2024) to construct a human animation framework. DynaScene (Yao et al., 2025) further introduces a multi-task learning approach to enhance the realism of the generated background . In comparison to the existing methods, our AnimateAnywhere predicts the background motion from human pose sequences without the reference video and camera pose extraction during inference.
2.2. Camera Control Video Generation
@@
With the impressive capabilities of video diffusion models (Ho et al., 2022; Guo et al., 2023; Blattmann et al., 2023; Bar-Tal et al., 2024; Gupta et al., 2025) in generating high-quality videos, several approaches have focused on achieving controllable video generation by incorporating conditional controls (Zhang et al., 2023b; Guo et al., 2025; Wei et al., 2024; Guo et al., 2023; He et al., 2024). In this section, we provide a brief overview of the progress in camera-controllable video generation. AnimateDiff (Guo et al., 2023) uses MotionLoRA (Hu et al., 2021) fine-tuning to adapt a pre-trained video model to specific camera movements. Direct-a-video (Yang et al., 2024a) introduces a camera embedder that incorporates three camera parameters, though the limited parameters restrict camera movement to translations and zooms. Following this, MotionCtrl (Wang et al., 2024c) incorporates rotation and translation matrices derived from extrinsic camera parameters, enabling the generation of videos with more complex camera motions. CameraCtrl (He et al., 2024) leverages Plücker coordinates to represent camera poses, allowing for geometric interpretations at the pixel level. Recent advances further enhance the geometric consistency of generated videos by epipolar constraints or 3D reconstruction. CamCo (Xu et al., 2024a) applies epipolar constraints to ensure consistency between generated frames and a single reference frame only. CAMI2V (Zheng et al., 2024) extends epipolar constraints onto all frames to account for scenarios where frames lack overlapping regions with the reference frame. CamTrol (Hou et al., 2024) models camera movements through explicit 3D point clouds. Building on the epipolar constraint used in CAMI2V (Zheng et al., 2024), we introduce an adaptive epipolar constraint on the 3D attention map of the background regions.
3. Method
@@
Given a human image , a background image , and a human pose sequence , AnimateAnywhere generates a video with synchronized human and background motion, guided solely by the pose sequence without explicit background motion control. Firstly, the preliminaries of AnimateAnywhere, including our base model CogVideoX (Yang et al., 2024b) and 3D Full Attention, are introduced in section 3.1. Additionally, we introduce the Background Motion Learner (BML) module that infers background motion representations from the latent features of the human pose sequence (as detailed in section 3.2). To further enhance the accuracy of background motion prediction, we incorporate epipolar masks derived from labeled camera trajectories as supervision (see section 3.3). Finally, the training objectives for AnimateAnywhere are described in section 3.4.
3.1. Preliminaries
@@
Image-to-video CogVideoX Model. CogVideoX (Yang et al., 2024b) is a video diffusion model with transformers (DiTs) that achieves high-quality video generation. Our AnimateAnyWhere is built upon the Image-to-Video framework of CogVideoX (Yang et al., 2024b). Given a condition image, CogVideoX (Yang et al., 2024b) first pads it with zeros to N frames. Then the padded input is encoded by VAE and concatenated with noise. Subsequently, the combined latent is iteratively denoised over several steps to generate the final video. CogVideoX (Yang et al., 2024b) employs unified 3D full attention modules within each DiT block. The 3D full attention module performs calculations simultaneously in spatial and temporal dimensions, capturing both intra-frame and cross-frame interactions. This enables CogVideoX (Yang et al., 2024b) to comprehensively aggregate temporal information for video generation.
@@ 3.2. AnimateAnywhere
Human Animation. (1) Reference Human Input. To ensure that the human details are well preserved during animation, we introduce a ReferenceNet DiT, whose structure matches the first 18 blocks of the denoising DiT. Specifically, the reference human image is first encoded into latent space using a 3D VAE and subsequently passed through the ReferenceNet DiT to extract human features . The human features are concatenated with the features from the main denoising DiT. These combined features are then processed through 3D full attention layers while preserving the original dimensions of the denoising DiT. (2) Human Pose Sequence Guidance. We adopt an architecture similar to ControlNet (Zhang et al., 2023a; Mou et al., 2024) to guide human motion. Specifically, we replicate the first 18 blocks of the pre-trained denoising DiT to serve as the ControlNet DiT. We utilize the VAE encoder to transform the human pose sequence into latent features and then utilize the ControlNet DiT to extract multi-scale pose features . Subsequently, these multi-scale features are added to the corresponding features of the denoising DiT via zero-initialized linear layers.
Background Animation. (1) Reference Background Input. Given a condition image, the CogVideoX (Yang et al., 2024b) I2V model uses it as the first frame to generate a video. Following this setup, we treat the reference background as the condition image and the initial background of the generated video. The background image is initially padded to match the dimension of the video, then encoded into latent features via the VAE encoder, and finally concatenated with noise to serve as input for the DiT module. (2) Background Motion Learner Module. Unlike other methods that extract camera trajectories from reference videos as background motion signals, we propose to directly predict background motion from human pose sequences. An intuitive approach is to learn camera trajectories from human pose sequences and incorporate the learned camera trajectories during video generation. However, directly learning camera trajectories faces several challenges. First, a single human pose may correspond to multiple plausible camera trajectories due to the ambiguity of whether the human is moving or stationary. Moreover, the labeled camera trajectories often suffer from scale inconsistency, which also increases the optimization difficulty. In contrast, rather than explicitly decoupling camera trajectories, we directly utilize the background motion features embedded in human pose to drive background movement. In particular, we introduce a LoRA-based module named Background Motion Learner (BML) module into each DiT block, and this module follows the standard paradigm of LoRA (Hu et al., 2022), expressed as . We add the human pose features and the denoising features together to predict the background motion. The process is as follows:
@@
3.3. Epipolar Constraint on 3D Attention Map
@@
Epipolar Line and Epipolar Mask. The camera poses consist of intrinsic parameters and extrinsic parameters , where represents the rotational matrix and ddenotes the translational matrix. The epipolar constraint establishes that a point in the -th frame must lie on its corresponding epipolar line in the -th frame, as illustrated in fig. 3. The epipolar line can be determined by camera poses as:
@@
where denotes a statistical threshold corresponding to the 30th percentile of attention values. Meanwhile, we emphasize that our loss function is equivalent to computing the difference between and the target map, as illustrated in fig. 2 (b)-(5). During inference, our AnimateAnywhere no longer requires epipolar information or camera trajectories, and it directly leverages the attention learned through the epipolar loss.
3.4. Learning Objective
We employ the latent loss in CogVideoX (Yang et al., 2024b) to train our pipeline,
@@ 3.4. Learning Objective
where and are the weights to balance the loss terms.
4. Experiments
4.1. Experimental Details
Dataset. We collect 5K videos from the Humanvid (Wang et al., 2024a) dataset and 3K videos from the Internet for training, where the Humanvid (Wang et al., 2024a) dataset has the annotation of camera pose sequence. Our training dataset consists of videos captured under diverse camera conditions, including both static and moving cameras.
We evaluate our model on the Humanvid (Wang et al., 2024a) test dataset and a wild dataset, Bilibili200 (BL200), collected from the internet. The Humanvid test dataset includes 100 videos with noticeable background motion but limited human motion. The BL200 dataset contains 200 videos with significant motion in both the human and background.
@@ 4.1. Experimental Details
Evaluation Metrics. Following the metrics utilized in (Hu, 2024; Xu et al., 2024b; Zhu et al., 2024; Zhou et al., 2024; Wang et al., 2024e, d), we use PSNR (Huynh-Thu and Ghanbari, 2008), SSIM (Wang et al., 2004), and LIPIS (Zhang et al., 2018) metrics to evaluate the quality of single-frame images. Additionally, we introduce FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019) metrics to evaluate the difference between the distributions of generated videos and real-world videos.
Implementation Details. We adopt the pretrained CogVideoX1.5-5B-I2V (Yang et al., 2024b) as our base model and keep it frozen throughout the training process. Both ControlNet and ReferenceNet are initialized with the parameters of CogVideoX (Yang et al., 2024b). Our training consists of three stages. In the first stage, we incorporate ControlNet into the base model and train only ControlNet while keeping all other parameters fixed. In the second stage, we introduce ReferenceNet and train it exclusively. In the final stage, we jointly optimize ControlNet, ReferenceNet, and the BML module.
In terms of data preparation, we first employ Grounded-SAM (Ren et al., 2024) to extract mask sequences in the human body and hand regions for each video. Subsequently, we obtain the human sequence and the background sequence using the extracted mask sequence. We randomly select a frame in the human sequence as the reference human and take the first frame from the background sequence as the reference background. Specifically, we complete the reference background using the inpainting method (Zhou et al., 2023). All experiments are trained on a mixture of horizontal (1360768) and vertical (7681360) videos and conducted on two A100 GPUs. Training is performed over 20,000 steps with a batch size of 8. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) with a constant learning rate of . We set the trade-off weights and to 0.2 and 0.005, respectively. During inference, classifier-free guidance (CFG) is applied with a scale of 2.5 for both reference images and pose sequences.
@@
PSNR | SSIM | LIPIS | FID | FVD | PSNR | SSIM | LIPIS | FID | FVD | |
MagicAnimate (Xu et al., 2024b) | 16.01 | 0.5463 | 0.4455 | 78.01 | 1943.07 | 16.82 | 0.5914 | 0.4133 | 40.02 | 1884.48 |
AnimateAnyone (Hu, 2024) | 16.47 | 0.5311 | 0.3913 | 50.18 | 1431.43 | 17.58 | 0.5756 | 0.3635 | 24.46 | 1507.05 |
Champ (Zhu et al., 2024) | 15.08 | 0.5075 | 0.4481 | 63.92 | 1837.07 | 16.88 | 0.5649 | 0.3919 | 30.69 | 1701.45 |
MusePose (Tong et al., 2024) | 16.93 | 0.5578 | 0.3676 | 52.00 | 1784.85 | 18.67 | 0.6126 | 0.3220 | 22.86 | 1580.47 |
AnimateAnyone† (Hu, 2024) | 16.81 | 0.5486 | 0.3672 | 47.69 | 1235.65 | 18.23 | 0.5930 | 0.3379 | 23.27 | 1371.22 |
MusePose† (Tong et al., 2024) | 17.33 | 0.5629 | 0.3563 | 44.73 | 1362.14 | 18.60 | 0.6116 | 0.3245 | 22.09 | 1444.79 |
Humanvid‡ (Wang et al., 2024a) | 16.42 | 0.5242 | 0.3676 | 47.41 | 1114.55 | - | - | - | - | - |
@@
4.2. Comparison with Existing Methods

### v2

@@
Abstract
Human image animation aims to generate human videos of given characters and backgrounds that adhere to the desired pose sequences. However, existing methods focus more on human actions while neglecting the generation of background, which typically leads to static results or inharmonious movements. The community has explored camera pose-guided animation tasks, yet preparing the camera trajectory is impractical for most entertainment applications and ordinary users. As a remedy, we present an AnimateAnywhere framework, rousing the background in human image animation without requirements on camera trajectories. In particular, based on our key insight that the movement of the human body often reflects the motion of the background, we introduce a background motion learner (BML) to learn background motions from human pose sequences. To encourage the model to learn more accurate cross-frame correspondences of background, we further deploy an epipolar constraint on the 3D attention map. Specifically, the mask used to suppress geometrically unreasonable attention is carefully constructed by combining an epipolar mask and the current 3D attention map. Extensive experiments demonstrate that our AnimateAnywhere effectively learns the background motion from human pose sequences, achieving state-of-the-art performance in generating human animation results with vivid and realistic backgrounds. Project page: https://animateanywhere.github.io/
Index Terms:
Video generation, human animation, background motion, diffusion model.
I Introduction
@@
Recent advancements in generative diffusion models have significantly advanced the field of image animation [1, 2, 3, 6, 7, 8], enabling the creation of highly realistic and controllable human image animation, particularly in terms of identity preservation and motion consistency. Despite these advancements, most existing human image animation methods focus on improving the generation quality of foreground humans while leaving the background static or unconstrained. However, in practice, the photographers usually need to adjust the cameras to better capture human movements, postures, and etc., which results in background changes. Thus, the static or disharmonious backgrounds are insufficient to meet the requirements for realistic and diverse human video generation.
To address this issue, one possible solution is to employ control signals to guide background motion. For example, Humanvid [4] controls the background motion via camera trajectories estimated from the reference videos, while Liu et al. [5] leverage sparse tracking points and human pose sequences to independently control background and foreground motions. However, these methods typically acquire camera trajectories and human pose sequences from a reference video, which is unavailable in many scenarios, such as science fiction film production. Although camera trajectories can be manually designed, this easily results in disharmony between human and background motion. It requires professional skills and is redundant for most applications and ordinary users. Additionally, another limitation of these methods is that a reference image with both the desired human (e.g., a clown) and desired background (e.g., the user’s living room) is required.
@@
Specifically, we build the AnimateAnywhere framework upon a pre-trained image-to-video model (i.e., CogVideoX [9]) and independently inject a reference human image, a reference background image, and a human pose sequence into the generation process. Thus, AnimateAnywhere supports free combinations of reference humans, backgrounds, and pose sequences, enabling highly flexible and controllable video generation. To realize the key idea of AnimateAnywhere, we propose the Background Motion Learner (BML) module, which learns background motion from the latent features of human pose sequences by applying LoRA [10] on the 3D-attention-based block. Moreover, although 3D full attention allows each background pixel to attend to any other pixel from all frames, this highly free attention manner also brings difficulties to optimization, easily resulting in geometrically inconsistent background generation. To further improve the cross-frame geometric consistency of the background, we propose an adaptive epipolar constraint for the 3D attention map. Therein, a mask used for suppressing geometrically unreasonable attention is elaborately designed from an epipolar mask and the current 3D attention map.
@@
We propose a Background Motion Learner module to predict background motion from human pose sequences and an adaptive epipolar constraint for the 3D attention map to enhance background geometric consistency.
@@
II Related Work
II-A Human Image Animation
@@
Human image Animation aims to generate coherent human videos from a reference image and a target pose sequence. The GAN-based methods [11, 12, 13, 14, 15, 16, 17] typically use deformation functions to spatially transform the reference image according to the input motion, thereby generating continuous video frames. However, these methods often suffer from unrealistic visual artifacts and temporal inconsistencies in the generated video. By leveraging the generative visual capabilities of diffusion models, recent advancements [1, 2, 3, 6, 7, 8, 4, 18] have enabled the generation of high-quality and consistent human animation videos. AnimateAnyone [1] and MagicAnimate [2] introduce a video diffusion model and a reference network for temporal consistency and identity preservation, respectively. Champ [3] incorporates four distinct control signals to provide comprehensive 3D shape and detailed pose guidance. Following these, some works focus on realistic hands [6, 7], efficient temporal modeling [8], or a pose alignment algorithm [19] to further improve the performance of human animation. Animate-X [20] extends animation to diverse character types with a unified framework.
@@
Despite achieving reasonable results in maintaining temporal consistency in humans, these approaches face challenges in generating realistic and dynamic backgrounds. MIMO [21] and AnimateAnyone2 [22] address this challenge by taking the background video directly from a reference video as input. Liu et al. [5] extract sparse point tracking from the reference video and use it to guide background animation. Humanvid [4] utilizes a SLAM-based method [23, 24] to estimate an accurate camera trajectory from the original videos and leverages CameraCtrl [25] and Animate Anyone [1] to construct a human animation framework. DynaScene [18] further introduces a multi-task learning approach to enhance the realism of the generated background. In comparison to the existing methods, our AnimateAnywhere predicts the background motion from human pose sequences without the reference video and camera pose extraction during inference.
II-B Camera Control Video Generation
@@
With the impressive capabilities of video diffusion models [26, 27, 28, 29, 30] in generating high-quality videos, several approaches have focused on achieving controllable video generation by incorporating conditional controls [31, 32, 33, 27, 25, 34]. In this section, we provide a brief overview of the progress in camera-controllable video generation. AnimateDiff [27] uses MotionLoRA [35] fine-tuning to adapt a pre-trained video model to specific camera movements. Direct-a-video [36] introduces a camera embedder that incorporates three camera parameters, though the limited parameters restrict camera movement to translations and zooms. Following this, MotionCtrl [37] incorporates rotation and translation matrices derived from extrinsic camera parameters, enabling the generation of videos with more complex camera motions. CameraCtrl [25] leverages Plücker coordinates to represent camera poses, allowing for geometric interpretations at the pixel level. Recent advances further enhance the geometric consistency of generated videos by epipolar constraints or 3D reconstruction. CamCo [38] applies epipolar constraints to ensure consistency between generated frames and a single reference frame only. CAMI2V [39] extends epipolar constraints onto all frames to account for scenarios where frames lack overlapping regions with the reference frame. CamTrol [40] models camera movements through explicit 3D point clouds. Building on the epipolar constraint used in CAMI2V [39], we introduce an adaptive epipolar constraint on the 3D attention map of the background regions.
III Method
@@
Given a human image , a background image , and a human pose sequence , AnimateAnywhere generates a video with synchronized human and background motion, guided solely by the pose sequence without explicit background motion control. Firstly, the preliminaries of AnimateAnywhere, including our base model CogVideoX [9] and 3D Full Attention, are introduced in section III-A. Additionally, we introduce the Background Motion Learner (BML) module that infers background motion representations from the latent features of the human pose sequence (as detailed in section III-B). To further enhance the accuracy of background motion prediction, we incorporate epipolar masks derived from labeled camera trajectories as supervision (see section III-C). Finally, the training objectives for AnimateAnywhere are described in section III-D.
III-A Preliminaries
@@
Image-to-video CogVideoX Model. CogVideoX [9] is a video diffusion model with transformers (DiTs) that achieves high-quality video generation. Our AnimateAnywhere is built upon the Image-to-Video framework of CogVideoX [9]. Given a condition image, CogVideoX [9] first pads it with zeros to N frames. Then the padded input is encoded by the VAE and concatenated with noise. Subsequently, the combined latent is iteratively denoised over several steps to generate the final video. CogVideoX [9] employs unified 3D full attention modules within each DiT block. The 3D full attention module performs calculations simultaneously in spatial and temporal dimensions, capturing both intra-frame and cross-frame interactions. This enables CogVideoX [9] to comprehensively aggregate temporal information for video generation.
@@ III-B AnimateAnywhere
Human Animation. (1) Reference Human Input. To ensure that the human details are well preserved during animation, we introduce a ReferenceNet DiT, whose structure matches the first 18 blocks of the denoising DiT. Specifically, the reference human image is first encoded into latent space using a 3D VAE and subsequently passed through the ReferenceNet DiT to extract human features . The human features are concatenated with the features from the main denoising DiT. These combined features are then processed through 3D full attention layers while preserving the original dimensions of the denoising DiT. (2) Human Pose Sequence Guidance. We adopt an architecture similar to ControlNet [41, 42] to guide human motion. Specifically, we replicate the first 18 blocks of the pre-trained denoising DiT to serve as the ControlNet DiT. We utilize the VAE encoder to transform the human pose sequence into latent features and then utilize the ControlNet DiT to extract multi-scale pose features . Subsequently, these multi-scale features are added to the corresponding features of the denoising DiT via zero-initialized linear layers.
Background Animation. (1) Reference Background Input. Given a condition image, the CogVideoX [9] I2V model uses it as the first frame to generate a video. Following this setup, we treat the reference background as the condition image and the initial background of the generated video. The background image is initially padded to match the dimensions of the video, then encoded into latent features via the VAE encoder, and finally concatenated with noise to serve as input for the DiT module. (2) Background Motion Learner Module. Unlike other methods that extract camera trajectories from reference videos as background motion signals, we propose to directly predict background motion from human pose sequences. An intuitive approach is to explicitly learn camera trajectories from human pose sequences and incorporate the learned camera trajectories during video generation. However, directly learning camera trajectories faces several challenges. First, splitting the task into two stages leads to error accumulation. Estimating camera trajectories in the form of RT matrices is inherently challenging, and any prediction error will be directly propagated and fixed in the second stage, which is difficult to correct even with video generation priors. Moreover, the labeled camera trajectories often suffer from scale inconsistency, which also increases the optimization difficulty. In contrast, rather than explicitly decoupling camera trajectories, we directly utilize the background motion features embedded in human pose to drive background movement. In particular, we introduce a LoRA-based module named Background Motion Learner (BML) module into each DiT block, and this module follows the standard paradigm of LoRA [10], expressed as . We add the human pose features and the denoising features together to predict the background motion. The process is as follows:
@@
III-C Epipolar Constraint on 3D Attention Map
@@
Epipolar Line and Epipolar Mask. The camera poses consist of intrinsic parameters and extrinsic parameters , where represents the rotational matrix and denotes the translational matrix. The epipolar constraint establishes that a point in the -th frame must lie on its corresponding epipolar line in the -th frame, as illustrated in fig. 3. The epipolar line can be determined by camera poses as:
@@
where denotes a statistical threshold corresponding to the 30th percentile of attention values. Meanwhile, we emphasize that our loss function is equivalent to computing the difference between and the target map, as illustrated in fig. 2 (b)-(5). During inference, our method no longer requires epipolar information or camera trajectories, and it directly leverages the attention learned through the epipolar loss.
III-D Learning Objective
We employ the latent loss in CogVideoX [9] to train our pipeline,
@@ III-D Learning Objective
where and are the weights to balance loss terms.
IV Experiments
IV-A Experimental Details
Dataset. We collect 5K videos from the Humanvid [4] dataset and 3K videos from the Internet for training, where the Humanvid [4] dataset has the annotation of camera pose sequence. Our training dataset consists of videos captured under diverse camera conditions, including both static and moving cameras.
We evaluate our model on the Humanvid [4] test dataset and a wild dataset, Bilibili200 (BL200), collected from the internet. The Humanvid test dataset includes 100 videos with noticeable background motion but limited human motion. The BL200 dataset contains 200 videos with significant motion in both the human and background.
@@ IV-A Experimental Details
Evaluation Metrics. Following the metrics utilized in [1, 2, 3, 6, 7, 8], we use PSNR [43], SSIM [44], and LPIPS [45] metrics to evaluate the quality of single-frame images. Additionally, we introduce FID [46] and FVD [47] metrics to evaluate the difference between the distributions of generated videos and real-world videos.
Implementation Details. We adopt the pretrained CogVideoX1.5-5B-I2V [9] as our base model and keep it frozen throughout the training process. Both ControlNet and ReferenceNet are initialized with the parameters of CogVideoX [9]. Our training consists of three stages. In the first stage, we incorporate ControlNet into the base model and train only ControlNet while keeping all other parameters fixed. In the second stage, we introduce ReferenceNet and train it exclusively. In the final stage, we jointly optimize ControlNet, ReferenceNet, and the BML module.
In terms of data preparation, we first employ Grounded-SAM [48] to extract mask sequences in the human body and hand regions for each video. Subsequently, we obtain the human sequence and the background sequence using the extracted mask sequence. We randomly select a frame in the human sequence as the reference human and take the first frame from the background sequence as the reference background. Specifically, we complete the reference background using the inpainting method [49]. All experiments are trained on a mixture of horizontal (1360768) and vertical (7681360) videos and conducted on two A100 GPUs. Training is performed over 20,000 steps with a batch size of 8. We employ the AdamW optimizer [50] with a constant learning rate of . We set the trade-off weights and to 0.2 and 0.005, respectively. During inference, classifier-free guidance (CFG) is applied with a scale of 2.5 for both reference images and pose sequences.
@@
| PSNR | SSIM | LPIPS | FID | FVD | PSNR | SSIM | LPIPS | FID | FVD | |
| MagicAnimate [2] | 16.01 | 0.5463 | 0.4455 | 78.01 | 1943.07 | 16.82 | 0.5914 | 0.4133 | 40.02 | 1884.48 |
| AnimateAnyone [1] | 16.47 | 0.5311 | 0.3913 | 50.18 | 1431.43 | 17.58 | 0.5756 | 0.3635 | 24.46 | 1507.05 |
| Champ [3] | 15.08 | 0.5075 | 0.4481 | 63.92 | 1837.07 | 16.88 | 0.5649 | 0.3919 | 30.69 | 1701.45 |
| MusePose [19] | 16.93 | 0.5578 | 0.3676 | 52.00 | 1784.85 | 18.67 | 0.6126 | 0.3220 | 22.86 | 1580.47 |
| DynamiCtrl [51] | 15.15 | 0.5054 | 0.4032 | 46.16 | 1907.71 | 16.22 | 0.5753 | 0.3511 | 26.75 | 2025.79 |
| AnimateAnyone† [1] | 16.81 | 0.5486 | 0.3672 | 47.69 | 1235.65 | 18.23 | 0.5930 | 0.3379 | 23.27 | 1371.22 |
| MusePose† [19] | 17.33 | 0.5629 | 0.3563 | 44.73 | 1362.14 | 18.60 | 0.6116 | 0.3245 | 22.09 | 1444.79 |
| DynamiCtrl† [51] | 17.27 | 0.5694 | 0.3462 | 44.84 | 1205.63 | 18.19 | 0.5901 | 0.3340 | 22.70 | 1565.75 |
| Humanvid∗ [4] | 17.91 | 0.5774 | 0.3007 | 46.66 | 991.66 | - | - | - | - | - |
@@
IV-B Comparison with Existing Methods